OpenStudy (anonymous):

The sum of three consecutive numbers is 48. What are the smallest of these numbers

OpenStudy (johnweldon1993):

so first number =x second = x + 1 and third = x + 2 x + x+1 + x+2 = 48 combine all your x values and combine all the constants and solve for 'x'

OpenStudy (anonymous):

x+x+1+x+2=48. 3x+3=48. Subtract 3 and get 3x=45. Divide both sides by three and get 15.

OpenStudy (anonymous):

Thanks